OpenStudy (anonymous):

Suppose a triangle has two sides of length 42 and 35, and that the angle between these two sides is 120. What is the length of the third side of the triangle?

OpenStudy (mathmale):

@hanzi: are you sure about using Heron's formula here? Isn't that formula for finding AREA? Want to think this through and perhaps propose a proper method of finding the length of the third side of the triangle?

OpenStudy (anonymous):

sorry, just slipped my mind, its law of sines

OpenStudy (anonymous):

ok so what the formula?

OpenStudy (anonymous):

formula is same, just change of name

OpenStudy (anonymous):

use law of sines formula, sina/A=sinb/B=sinc/C A=42, B=35, c=120 put in above equation, also use a+b+c=180 (sum of angles of triangles) youll get value of C, the third side

OpenStudy (mathmale):

@Hanzi: I'm afraid that 's not true. If you believe the Law of Sines applies here, demonstrate it. There is another "Law" that is more relevant to this problem.

OpenStudy (anonymous):

what would be the other Law that i could use?

OpenStudy (mathmale):

Please look up "Law of Cosines." Once you've read about that and copied down the formula, I'd be happy to help you apply it to this particular homework problem.

OpenStudy (anonymous):

ok i will reply back when i have.

OpenStudy (mathmale):

I like your attitude. Go for it!

OpenStudy (anonymous):

c2=a2+b2-2ab cos(C)? is that right?

OpenStudy (anonymous):

would the answer be 66.78?

OpenStudy (mathmale):

You have the correct formula. Just checking: What does C stand for in this case? c? What are the lengths of a and b?

OpenStudy (mathmale):

\[IF: ~c^2=a^2+b^2-2ab*\cos C,~then~here:~c^2=35^2+42^2-2(35)(42)\cos(120)\]

OpenStudy (mathmale):

Please calculate this.